OpenStudy (anonymous):

Evaluate -3x3 - 4x for x = -1.

OpenStudy (anonymous):

-5

OpenStudy (anonymous):

how fid you get that

OpenStudy (anonymous):

*did

OpenStudy (anonymous):

-3x3 =-9 -4x-1 = 4 -9+4 = -5

OpenStudy (anonymous):

thank you